A boy on a bridge throws a stone horizontally with a speed of 25 m/s, releasing the stone from a point 19.6 m above the surface of the river. How far from a point directly below the boy will the stone strike the water?


1
Expert's answer
2021-09-15T11:39:34-0400

Given:

"v_0=25\:\rm m/s"

"H=19.6\:\rm m"


The horizontal distance

"L=v_0\sqrt{\frac{2H}{g}}=25\sqrt{\frac{2*19.6}{9.8}}=50\:\rm m"